# SensorChat: Answering Qualitative and Quantitative Questions during Long-Term Multimodal Sensor Interactions

### v1

@@ 6.3. Qualitative Performance
Sensor+Text IMU2CLIP-GPT4 (Moon et al., 2023b) FSL 0.44 0.28 0.40 0.13 0.18
@@
Table 4. Quantitative results of SensorChat compared against state-of-the-art methods. Bold and underlined values show the best results (all achieved by SensorChat) and the best among baselines.
6.4. Quantitative Performance
@@
Table 4 presents the quantitative results of all methods on the three variants of SensorQA (Reichman et al., 2025): full answers, short answers, and multiple-choice. Note that DeepSQA (Xing et al., 2021) was not evaluated on the multiple-choice version due to its inability to handle dynamic answer choices. It is important to highlight that exact match accuracy for short and multiple-choice answers is a strict metric, as it requires the model to generate answers in the exact same form as the correct ones. For instance, ”4 hours” and ”3 hours 50 min” would be considered different. Answering multiple-choice questions can also be particularly challenging, as candidate answers differ only slightly, such as ”A. 10 min” vs. ”B. 20 min,” making them difficult to distinguish. In practical applications, however, a QA system does not necessarily need to achieve perfect exact match accuracy to be useful. We leave the exploration of more advanced metrics that better align with user satisfaction for future work, as discussed in Sec. 8.
As shown in Table 4, our SensorChat outperforms the best state-of-the-art methods with the highest Rouge scores on full answers, 26% higher accuracy on short answers, and 14% higher accuracy on multiple choices. The top Rouge scores highlight SensorChat’s ability to generate high-quality natural language that are the most similar with the ground-truth answers in SensorQA (Reichman et al., 2025). Even under the strict exact match evaluation, SensorChat achieves 54% accuracy on short answers and 70% on multiple-choice questions. These accuracy improvements demonstrate SensorChat’s effectiveness in learning the underlying activity facts from the long-duration, multimodal time series sensor data. As explained in Sec. 6.3, all three stages are crucial for achieving high accuracy, including correct question decomposition to invoke the right functions, precise sensor queries to accurately extract activity information from sensor data, and effective answer assembly for generating natural language responses. SensorChatC performs slightly better than SensorChatE mainly due to the stronger model capability of GPTs compared to quantized LLaMA in question decomposition. However, SensorChatE, as a pure edge solution, preserves better user privacy as discussed in Sec. 5.
In contrast, all baselines struggle with quantitative accuracy, with an highest accuracy of merely 28% on the short answers. Despite GPT-4’s strong reasoning capabilities and its multimodal variants (GPT-4-Turbo and GPT-4o), the models do not perform optimally for processing sensor data and daily life activities, as they are not specifically trained for these tasks. The Rouge scores are lower because GPTs tend to generate longer text, resulting in less similarity with the ground-truth answers. Interestingly, text-only baselines with finetuning, such as LLaMA2-7B (Touvron et al., 2023a) and T5 (Raffel et al., 2020), achieve some of the highest accuracy among the baselines, with 27% accuracy on short answers and 56% accuracy on multiple-choice datasets. These results suggest a sensor bias in the dataset, where some questions can be “guessed” correctly without sensor data. Accuracy scores lower than these baselines indicate ineffective fusion of text and sensor data, as seen with models like Llava-1.5 (Liu et al., 2024a) and OneLLM (Han et al., 2024), whose poor performance stems from mismatches in pretraining and finetuning data formats. For example, OneLLM was designed for aligning raw IMU signals with fixed window sizes, making it challenging to adapt to long-duration sensor data in our task. Finally, DeepSQA (Xing et al., 2021) and LLaMA-Adapter (Zhang et al., 2023) perform best among the baselines but struggle with accurate quantitative analysis on long-duration time series, as discussed in Sec.3 and Sec.6.3.
@@
Figure 9. The answer accuracy and latency trade-offs of all methods measured on the cloud and edge platforms. We evaluate SensorChatC on A100 (a10, 2025) and SensorChatE on Jetson Orin NX (jet, 2025). All LLM-based models are quantized to 4-bit weights with AWQ (Lin et al., 2024a) for Jetson Orin deployments.
@@ 6.5. End-to-End Answer Generation Latency
To evaluate efficiency, we measure the end-to-end answer generation latency of SensorChatC and SensorChatE on their respective platforms - cloud server and NVIDIA Jetson Orin. For a fair comparison, we quantize all LLM-based baselines to 4-bit weights using AWQ (Lin et al., 2024a) for the Jetson Orin experiments, matching SensorChatE ’s setup. SensorChat takes an average generation latency of 2.3s on the cloud and 10.0s on the Jetson Orin. Specifically, SensorChatC takes an average of 1.2s for question decomposition, 0.5s for sensor data query, and 0.6s for answer assembly. SensorChatE takes an average of 2.5s, 4.9s, and 2.5s for each stage, respectively.

### v2

@@ 6.3. QA Performance
Sensor+Text IMU2CLIP+GPT4o (Moon et al., 2023b) FSL 0.58 0.44 0.55 0.28 0.49
Sensor+Text LLMSense (Ouyang and Srivastava, 2024) FSL 0.37 0.26 0.35 0.23 0.49
@@
Table 5. Comprehensive results of SensorChat compared against state-of-the-art methods on SensorQA (Reichman et al., 2025). Bold and underlined values show the best results (all achieved by SensorChat) and the best among baselines. DeepSQA (Xing et al., 2021) was not evaluated on the multiple-choice version due to its inability to handle dynamic answer choices.
@@
Table 5 presents the comprehensive results of all methods on the three variants of SensorQA (Reichman et al., 2025): full answers, short answers, and multiple-choice. The full answers dataset evaluates overall language quality, while the short answers and multiple-choice evaluations focus on the model’s ability to learn underlying facts and numerical results rather than the language itself. Our SensorChat outperforms the best state-of-the-art methods with the highest Rouge scores on full answers, 93% higher accuracy on short answers, and 25% higher accuracy on multiple choices. SensorChatC performs slightly better than SensorChatE mainly due to the stronger model capability of GPTs compared to quantized LLaMA in question decomposition. However, SensorChatE, as a pure edge solution, preserves better user privacy as discussed in Sec. 5. The gains of SensorChat over the baselines are more significant for short answers/multiple choices than for full answers, highlighting SensorChat’s particular advantage in factual and numerical reasoning in comparison to existing works. It is important to note that the exact match accuracy for short and multiple-choice answers is a strict metric, as it requires the model to generate the exact answer as the correct one. For instance, ”4 hours” and ”3 hours 50 min” would be considered different. Answering multiple-choice questions can also be particularly challenging, as candidate answers differ only slightly, such as ”A. 10 min” vs. ”B. 20 min,” making them difficult to distinguish. In practical applications, however, a QA system does not necessarily need to achieve perfect exact match accuracy to be useful. We leave the exploration of more advanced metrics that better align with user satisfaction for future work, as discussed in Sec. 8. Even under the strict exact match evaluation, SensorChat achieves 54% accuracy on short answers and 70% on multiple-choice questions. These accuracy improvements demonstrate SensorChat’s effectiveness in learning the underlying life facts from the long-duration, multimodal time series sensor data.
In contrast, all baseline methods struggle with quantitative accuracy, achieving at most only 28% on short-answer questions. Surprisingly, text-only baselines like LLaMA2-7B (Touvron et al., 2023a) and T5 (Raffel et al., 2020) achieve some of the highest accuracy among the baselines, with 27% accuracy on short answers. The limited improvements from adding vision or sensor modalities suggest that existing baselines fail to effectively fuse multimodal information in our setting. There are several reasons for this underperformance. GPT-based models (e.g., GPT-4-Turbo and GPT-4o (Achiam et al., 2023)) are not trained for interpreting sensor data or reasoning about daily activities. Their lower Rouge scores are often due to overly long answers that deviate from the ground truth. Vision+text models like LLaMA-Adapter (Zhang et al., 2023) and LLaVA-1.5 (Liu et al., 2024a) were originally developed for video captioning and lack the precision needed for quantitative reasoning in daily life activities. For sensor+text baselines, IMU2CLIP+GPT-4o (Moon et al., 2023b) and LLMSense (Ouyang and Srivastava, 2024) rely on feeding long narrative logs to LLMs to provide the sensor information necessary for accurately answering the questions, making them vulnerable to LLMs’ well-known struggles with long-context reasoning (Huang et al., 2023; Gu and Dao, 2023). Others like DeepSQA (Xing et al., 2021) and OneLLM (Han et al., 2024) were built for short-duration data (i.e., less than one minute) and cannot scale effectively to long-duration, high-frequency inputs of multiple days with their current model and training pipeline. In contrast, SensorChat overcomes these limitations by introducing a dedicated query stage that retrieves relevant sensor embeddings from a pre-encoded database. This design enables accurate answers to quantitative questions in SensorQA (Reichman et al., 2025).
@@
Figure 8. The answer accuracy and latency trade-offs of all methods measured on the cloud and edge platforms. We evaluate SensorChatC on A100 (a10, 2025) and SensorChatE on Jetson Orin NX (jet, 2025). All LLM-based models are quantized to 4-bit weights with AWQ (Lin et al., 2024) for Jetson Orin deployments.
@@ 6.4. End-to-End Answer Generation Latency
To evaluate efficiency, we measure the end-to-end answer generation latency of SensorChatC and SensorChatE on their respective platforms - cloud server and NVIDIA Jetson Orin. For a fair comparison, we quantize all LLM-based baselines to 4-bit weights using AWQ (Lin et al., 2024) for the Jetson Orin experiments, matching SensorChatE ’s setup. SensorChat takes an average generation latency of 2.3s on the cloud and 10.0s on the Jetson Orin. Specifically, SensorChatC takes an average of 1.2s for question decomposition, 0.5s for sensor data query, and 0.6s for answer assembly. SensorChatE takes an average of 2.5s, 4.9s, and 2.5s for each stage, respectively.